Find the non-trivial solution of the sturn_liouville problem d2y/dx2+λy=0 y(0)=0, y(π)=0


1
Expert's answer
2021-11-03T15:12:42-0400

Any eigenvalues of Equation "\dfrac{dy^2}{x^2}+λy=0" must be positive. If "y" satisfies the equation with "\lambda>0," then


"y=c_1\cos\sqrt{\lambda}x+c_2\sin\sqrt{\lambda}x"

where "c_1" and "c_2" are constants. The boundary condition "y(0)=0" implies that "c_1=0." Therefore


"y=c_2\sin\sqrt{\lambda}x"

The boundary condition "y(\pi) = 0" implies that


"c_2\sin\sqrt{\lambda}\pi=0"

To make "c_2\sin\sqrt{\lambda}\pi=0" with "c_2 \not= 0," we must choose "\sqrt{\lambda} = n," where "n" is a positive integer.

Therefore "\lambda_n^2=n^2" is an eigenvalue and


"y_n=\sin(nx)"

is an associated eigenfunction.